I'm having a difficult time with this problem could I get some assistance please?
Solve the system of equations by substitution:
Y=X^2+2X-2 AND y=3x+4

Answer by troyapplehelen(46)

y=X^2+2X-2
y=3X+4, so, X^2+2X-2 = 3X+4
then, x^2+2x-2-3x-4=0
x^2-x-2=o
(x-2)(x+1) = 0 ;t/f, x is 2 or -1.